Lets say a company has 10 members on its management team :7 men and 3 women. There is a conference in Hawaii, and the company has decided to send 3 people. If the selection of the 3 people to go to Hawaii is determined by lottery, what is the probability that at least 1 woman will be selected?
Answer by Boreal(15235)

probability of selecting at least 1 is 1-probability of not selecting any.
That would be (7/10)(6/9)(5/8)=210/720=0.292.
The probability of at least 1 is 0.708.